displacement x  ⟺  angular displacement θ

velocity v  ⟺  angular velocity ω

acceleration a  ⟺  angular acceleration α

v = \frac{dx}{dt}  ⟺  ω = \frac{dθ}{dt}

a = \frac{dv}{dt} = \frac{d^{2}x}{dt^{2}}  ⟺  α = \frac{dω}{dt} = \frac{d^{2}θ}{dt^{2}}

x = R θ

v = R ω

a = R α

x_{f} = x_{i} + vt + \frac{1}{2}at^{2}

x_{f} = x_{i} + \frac{1}{2}(v_{f} + v_{i})t

v_{f} = v_{i} + at

v_{f}^{2} = v_{i}^{2} + 2a(x_{f} − x_{i})

θ_{f} = θ_{i} + ωt + \frac{1}{2}αt^{2}

θ_{f} = θ_{i} + \frac{1}{2}(ω_{f} + ω_{i})t

ω_{f} = ω_{i} + αt

ω_{f}^{2} = ω_{i}^{2} + 2α(θ_{f} − θ_{i})